OpenStudy (anonymous):

the celsius temperature of a 3.00-L sample gas is lowered from 80.0 C to 30.0 C what will be the resulting volume of this gas

OpenStudy (chmvijay):

keeping pressure constant right u use PV=nRT equation

OpenStudy (anonymous):

right

OpenStudy (anonymous):

but i need to show steps:/

OpenStudy (chmvijay):

V1 / T1 = V2 / T2 can you now

OpenStudy (anonymous):

not really idk how to do it this

OpenStudy (chmvijay):

according to charles law this equation is V1=T1*V2 /T2 = 303 * 3 / 353 = 2.57 L

OpenStudy (chmvijay):

since temperature has benn reduced from 80 to 30 the volume of gas should decrease hence it decreased from 3 to 2.57 L

OpenStudy (chmvijay):

303 and 353 K are just i converted 30 and 80 °C to kelvin :)

OpenStudy (anonymous):

perfect